OpenStudy (anonymous):

Integrate ((cube root of x) - 1) / square root of x) with respect to x. Do I need to use substitution and what do I substitute?

OpenStudy (turingtest):

\[\int{\sqrt[3]x-1\over\sqrt x}dx\]if this is correct, rewrite this with fractional exponents and divide out the terms you can then integrate straightforward

OpenStudy (anonymous):

yes, it's correct. I will try and if I fail, I shall return for help

OpenStudy (turingtest):

I'll be around :)

OpenStudy (anonymous):

\[6\div5 \times \sqrt[6]{x ^{5}} + 2\sqrt{x}\] is this the answer, do you maybe know?

OpenStudy (turingtest):

sure, just don't forget the +C *always* include the +C for indefinite integrals

OpenStudy (anonymous):

oh yeah, thanks a lot :)

OpenStudy (turingtest):

np :)